Circular No. 8673

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5117 SAGITTARII = NOVA SAGITTARII 2006 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag ≈ 9.0) on three Technical Pan photographs taken with an 85mm-f.l. camera lens (+ orange filter) around Feb. 17.37 UT. The new object is located at α = 17^{h}58^{m}.9, δ = −36^{o}48′ (equinox 2000.0). Nothing brighter than mag 11.0 was seen at this position on a photo taken on Feb. 5.36. Liller adds that a low-dispersion CCD spectrogram of the Hα region (resolution 0.32 nm/pixel) of the apparent nova, taken on Feb. 19.38 UT with a 0.2-m Schmidt telescope (+ 75 lines/mm grating), shows a moderately strong Hα emission with a FWHM of 630 ± 25 km/s. A weak P-Cyg profile may be present ≈ 535 km/s from the peak of the Hα emission. H. Yamaoka, Kyushu University, reports the independent discovery of this nova at mag ∼ 8.6 on Feb. 17.84 by Minoru Yamamoto (Okazaki, Aichi, Japan) via two digital-camera images taken with a 85-mm-f.l. camera lens and a Fuji Finepix S2Pro camera; Yamamoto’s position for the new object is α = 17^{h}58^{m}54^{s}, δ = −36^{o}47′40′′. C. Jacques and E. Pimentel, Belo Horizonte, Brazil, write that their unfiltered CCD images taken on Feb. 19.34 UT show the new object at mag 8.6 with the following position end figures: 52^{s}.61, 35′′.1 (equinox 2000.0). Jacques adds that a USNO-B1.0 catalogued source that is also visible on a red 1979 Digitized Sky Survey plate (with limiting mag 19.9) has red mag 17.4 and position end figures 52^{s}.60, 36′′.2. Visual magnitude estimate by A. Pearce, Nedlands, W. Australia: Feb. 19.848 UT, 9.2. N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that the designation V5117 Sgr has been given to this nova.

RS OPHIUCHI R. K. Das, N. M. Ashok, and D. P. K. Banerjee, Physical Research

Laboratory, report that near-infrared JHK spectroscopy of RS Oph during its current outburst (cf. IAUC 8671), obtained with the Mt. Abu 1.2-m telescope (+ PRL Near Infrared NICMOS3 Imager/Spectrometer) on Feb. 16.99 UT, shows prominent H I emission lines of Paβ, Paγ, Brγ, Brackett 10–19, He I at 1.0830 and 2.0581 μm, and O I at 1.1288 μm.

CCD magnitudes by G. Sostero and E. Guido, Udine, Italy (via a remotely operated 0.25-m reflector near Mayhill, NM): Feb. 13.490 UT, B = 5.90; 13.491, V = 5.00; 13.493, R_{c} = 4.40; 14.487, B = 6.25; 14.488, V = 5.47; 14.490, R_{c} = 4.40; 16.484, B = 6.99; 16.485, V = 6.23; 16.487, R_{c} = 5.04; 20.481, B = 8.03; 20.483, V = 7.23; 20.485, R_{c} = 5.05.

2006 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green